a box is 3 times as long as it is wide.it is 6 inches long.how wide is the box
Answer by macston(5194)

L=Length=6 in, W=Width=(1/3)L
W=1/3(L)=1/3(6 in)=2 in
ANSWER The box is 2 inches wide.